OpenStudy (anonymous):

on a right angle if two sides are 110 and 100 what would x be. Pythagorean theorem.

OpenStudy (azureilai):

assuming that x is the hypotenuse 110^2+100^2=x^2

OpenStudy (mathstudent55):

Is there a figure?

OpenStudy (mathstudent55):

There are two possibilities: \(100^2 + 110^2 = x^2\) \(100^2 + x^2 = 110^2\)

OpenStudy (mathstudent55):

If the hypotenuse is x, then use the first equation. That means @Azureilai assumed correctly.